OpenStudy (anonymous):

In the given equation, for any real number X and Y, what is the value of a+b? -(Y-X)+2(3X+7)=by+ax+14

OpenStudy (campbell_st):

well looking at the right hand side of the equation, and with some distributing you get -y +x +6x + 14 collect like terms -y + 7x + 14 = by + ax + 14 so -y = by 7x = ax I'll let you finish

OpenStudy (campbell_st):

not quite -y really means \[-1 \times y\] and you are comparing \[-1 \times y = b \times y\] cross the y's out and what would b equal?

OpenStudy (campbell_st):

thats great now for a \[7 \times x = a \times x\] whats the value of a

OpenStudy (campbell_st):

thats it... well done